How fast must an object travel for its total energy to be 55% more than its rest energy? Answer in c.
Answer #1

E ( total enrgy) = \gamma E0

E0 =rest energy = mc^2

1.55 E0 = \gamma E0

\gamma = 1/ sqroot ( 1-v^2/ c^2)

1.55 sqroot ( 1- v^2/ c^2) = 1

( 1-v^2/ c^2) = 1/ 2.4025

0.584= v^2/ c^2

0.764= v/v

v = 0.764 c